solve by the substitution method.
4m + n= 2
m- 8n =50
What is the solution of the system?

Answer by mananth(16949)

4m+n=2
m=8n+50
substitute m
4(8n+50)+n=2
32n+200+n=2
33n=-198
/33
n=-6
plug n in m=8n+50
m=8(-6)+50
m=2